A blue laser beam with a wavelength of 450nm illuminates 2 narrow slits each with a width of 0.5micrometers. The opaque strip separating the strips is 15 micrometers wide. The interference pattern is observed on a screen 85cm away.

1. [2] Relative to the optical axis through the center of the slits, at what angle in radians does the 1st interference minima occur?

2. [2] How far from the axis is this (in cm)?

3. [2] At what angle does the 3rd interference maxima (m=3) occur?

4. [2] How far from the axis is this (in cm)?

5. [4] Sketch the pattern (as a function of angle) from 3 maxima below to 3 maxima above the axis.



1
Expert's answer
2021-03-09T15:28:37-0500

Given,

"\lambda = 450nm"

Width of the slit "(w) = 0.5\mu m"

Opaque strip separation "= 15\mu m"

Screen distance from the slit "(D)=85 cm=0.85m"

i) Angular distance "(\theta)=\sin^{-1}(\frac{\lambda}{(m+1)d})"

Here m = 1

"(\theta)=\sin^{-1}(\frac{\lambda}{(1+1)d})"

Now, substituting the values,

"(\theta)=\sin^{-1}(\frac{450\times 10^{-9}}{(2)\times (15+0.5)\times 10^{-6}})"

"=\sin^{-1}(0.0145)"

"\simeq 0.83 ^\circ"


ii) first interference minima "=\frac{D\lambda }{2d}"

Now, substituting the values,

"=\frac{0.5\times 0.85\times 450\times 10^{-9}}{2\times (15+0.5)}"


"=0.0123m"

"\simeq1.23cm"

iii)

For m=3,

"\theta = \sin^{-1}(\frac{m\lambda D}{d})"

Now, substituting the values,

"\theta = \sin^{-1}(\frac{3\times 450\times 10^{-9} m}{15.6\times 10^{-6}})"

"=\sin^{-1}(0.087)"

"=4.99^\circ"

"\simeq 5^\circ"

iv) here m = 3

"y=\frac{m \lambda D}{d}"

Now, substituting the values,

"y=\frac{3\times 450\times 10^{-9} m\times 0.85}{15.6\times 10^{-6}}"

"=0.074m"

"\simeq 7.4 cm"


v) The required pattern is given below-